# Dimension reduction and the gradient flow of relative entropy

### v1

@@ 1.1. The SNE and t-SNE algorithms
$p_{j|i}=\frac{\exp(-|x_{i}-x_{j}|^{2}/(2\sigma_{i}^{2}))}{\sum_{k\neq i}\exp(-|x_{i}-x_{k}|^{2}/(2\sigma_{i}^{2}))}\quad\textrm{for }i\neq j.$
This represents “the conditional probability that $x_{i}$ would pick $x_{j}$ as its neighbor if neighbors were picked in proportion to their probability density under a Gaussian centered at $x_{i}$[14], where $\sigma_{i}$ is the variance of the Gaussian. The numbers $\sigma_{i}$ are selected according to certain criteria; since this is not the focus of the paper refer the reader to [14]. We define
@@
(1.2) $\mathcal{C}(Y):=\sum_{i\neq j}p_{ij}\log\frac{p_{ij}}{q_{ij}},$
@@ 1.2. The gradient flow of relative entropy
(1.3) $\begin{split}\frac{dy_{i}}{dt}={}&4\sum_{j\neq i}(p_{ij}-q_{ij})(y_{i}-y_{j})(\log\beta)^{\prime}(|y_{i}-y_{j}|^{2}),\end{split}$
@@ 1.3. Results
    $\sup_{x\in\mathbb{R}}(\log\gamma)^{\prime}(x)<\infty.$
@@
There are some natural questions to ask regarding the long time behavior of $Y(t)$.
  1. (i)
  2. (ii)
  3. (iii)
  4. (iv)
@@ 1.4. SNE versus t-SNE
Our results show a marked difference in the behavior of the flow $Y(t)$ in the SNE versus t-SNE cases. Only t-SNE exhibits divergence of points at infinity, and our Theorem 1.3 shows that if this occurs one obtains distinct points in the rescaled limit. We believe this is related to the so-called “crowding problem” discussed by van der Maaten and Hinton in [14, Section 3.2]. They note that in SNE, the points $y_{i}$ tend to be pushed towards the center, preventing gaps from occurring between clusters. They give heuristic reasons why the heavier tails of the probability distribution in t-SNE may compensate for this effect, pushing points apart. Indeed, this motivated their construction of the t-SNE algorithm.
The outline of the paper is as follows. In Section 2 we compute the formula (1.3) for the gradient flow. Section 3 is the heart of the paper, in which we prove the main results, Theorems 1.1 to 1.4. Finally, in Section 4, we give two further examples.
Acknowledgements.  The author is very grateful to Antonio Auffinger for spurring his interest in the mathematical study of t-SNE, and for some helpful comments.
@@ 2. The gradient flow of relative entropy
In this section we prove the formula (1.3) for the gradient flow of the entropy $\mathcal{C}$. The gradient of $\mathcal{C}$ was calculated for $\beta=e^{-x}$ and $\beta(x)=(1+x)^{-1}$ in [7, 14] (see also [4, Chapter 16]) and its extension to general $\beta$ is straightforward. However, we include the details here for the convenience of the reader. Recalling (1.1), we write
(2.1) $q_{ij}=\frac{\beta(|y_{i}-y_{j}|^{2})}{Z},\quad\textrm{for }Z:=\sum_{k\neq\ell}\beta(|y_{k}-y_{\ell}|^{2}).$
Compute for $m\neq j$,
$\begin{split}\nabla_{y_{m}}q_{mj}={}&\frac{2}{Z^{2}}\left(Z\beta^{\prime}(|y_{m}-y_{j}|^{2})(y_{m}-y_{j})-2\beta(|y_{m}-y_{j}|^{2})\sum_{\ell\neq m}\beta^{\prime}(|y_{m}-y_{\ell}|^{2})(y_{m}-y_{\ell})\right).\end{split}$
On the other hand, if $m$ is not equal to either $i$ or $j$, we have
$\begin{split}\nabla_{y_{m}}q_{ij}={}&-\frac{4q_{ij}}{Z}\sum_{\ell\neq m}\beta^{\prime}(|y_{m}-y_{\ell}|^{2})(y_{m}-y_{\ell}).\end{split}$
Then
$\begin{split}\nabla_{y_{m}}\mathcal{C}={}&-\sum_{i\neq j}\frac{p_{ij}}{q_{ij}}\nabla_{y_{m}}q_{ij}\\ ={}&-2\sum_{j\neq m}\frac{p_{mj}}{q_{mj}}\nabla_{y_{m}}q_{mj}-\sum_{\begin{subarray}{c}i\neq j\\ i\neq m,j\neq m\end{subarray}}\frac{p_{ij}}{q_{ij}}\nabla_{y_{m}}q_{ij}\\ ={}&-4\sum_{j\neq m}\frac{p_{mj}}{Z\beta(|y_{m}-y_{j}|^{2})}\bigg{(}Z\beta^{\prime}(|y_{m}-y_{j}|^{2})(y_{m}-y_{j})\\ {}&-2\beta(|y_{m}-y_{j}|^{2})\sum_{\ell\neq m}\beta^{\prime}(|y_{m}-y_{\ell}|^{2})(y_{m}-y_{\ell})\bigg{)}\\ {}&+4\sum_{\begin{subarray}{c}i\neq j\\ i\neq m,j\neq m\end{subarray}}\frac{p_{ij}}{Z}\sum_{\ell\neq m}\beta^{\prime}(|y_{m}-y_{\ell}|^{2})(y_{m}-y_{\ell}).\end{split}$
But since $\sum_{i\neq j}p_{ij}=1$ we have
$\sum_{\begin{subarray}{c}i\neq j\\ i\neq m,j\neq m\end{subarray}}p_{ij}=1-2\sum_{j\neq m}p_{mj},$
and hence
$\begin{split}\nabla_{y_{m}}\mathcal{C}={}&-4\sum_{j\neq m}\frac{p_{mj}\beta^{\prime}(|y_{m}-y_{j}|^{2})}{\beta(|y_{m}-y_{j}|^{2})}(y_{m}-y_{j})\\ {}&+\frac{8}{Z}\left(\sum_{j\neq m}p_{mj}\right)\sum_{\ell\neq m}\beta^{\prime}(|y_{m}-y_{\ell}|^{2})(y_{m}-y_{\ell})\\ {}&+\frac{4}{Z}(1-2\sum_{j\neq m}p_{mj})\sum_{\ell\neq m}\beta^{\prime}(|y_{m}-y_{\ell}|^{2})(y_{m}-y_{\ell})\\ ={}&-4\sum_{\ell\neq m}\left(p_{m\ell}-q_{m\ell}\right)(y_{m}-y_{\ell})(\log\beta)^{\prime}(|y_{m}-y_{\ell}|^{2}),\\ \end{split}$
@@ 3. Proofs of the main results
$\mathcal{C}=\sum_{i\neq j}p_{ij}\log\frac{p_{ij}}{q_{ij}},$
we have for $i\neq j$,
$-p_{ij}\log q_{ij}\leq C,$
which implies in particular that
@@
for a uniform positive $c^{\prime}$. Recall that by assumption, each $p_{ij}$ is strictly positive.
@@
$\begin{split}\frac{d}{dt}\sum_{i=1}^{n}|y_{i}|^{2}={}&8\sum_{i\neq j}(p_{ij}-q_{ij})y_{i}\cdot(y_{i}-y_{j})(\log\beta)^{\prime}(|y_{i}-y_{j}|^{2})\\ ={}&8\sum_{i<j}(p_{ij}-q_{ij})y_{i}\cdot(y_{i}-y_{j})(\log\beta)^{\prime}(|y_{i}-y_{j}|^{2})\\ {}&+8\sum_{i<j}(p_{ij}-q_{ij})y_{j}\cdot(y_{j}-y_{i})(\log\beta)^{\prime}(|y_{i}-y_{j}|^{2})\\ ={}&8\sum_{i<j}(p_{ij}-q_{ij})|y_{i}-y_{j}|^{2}(\log\beta)^{\prime}(|y_{i}-y_{j}|^{2}).\end{split}$
Now for $i\neq j$, write
@@
$\begin{split}\frac{d}{dt}S={}&\frac{8}{Z}\left(p_{12}Z-A_{12}\right)\log\frac{A_{12}}{Z}+\frac{8}{Z}\sum_{\begin{subarray}{c}i<j\\ (i,j)\neq(1,2)\end{subarray}}\left(p_{ij}Z-A_{ij}\right)\log\frac{A_{ij}}{Z}\\ \leq{}&-\frac{4c^{\prime}}{Z}p_{12}ZS+\frac{8}{Z}\sum_{\begin{subarray}{c}i<j,(i,j)\neq(1,2)\\ p_{ij}Z-A_{ij}<0\end{subarray}}\left(p_{ij}Z-A_{ij}\right)\log\frac{A_{ij}}{Z},\end{split}$
@@
$\begin{split}\frac{d}{dt}S={}&-cS+\frac{8}{Z}\sum_{\begin{subarray}{c}i<j,(i,j)\neq(1,2)\\ p_{ij}Z-A_{ij}<0\end{subarray}}\left(p_{ij}Z-A_{ij}\right)\log p_{ij}\\ \leq{}&-cS+C<0,\end{split}$
@@
We provide an example with three points in $\mathbb{R}$. Write $p_{12}=p_{23}=a$ so that $p_{13}=1/2(1-4a)$ with $a\in(0,1/4)$ to be determined. Assume that $y_{1}(t)=X(t)$, $y_{2}(t)=0$ and $y_{3}(t)=-X(t)$ so that the points are symmetric about the origin, with $X(0)>0$. Recalling (2.1), calculate $Z=4\beta(X^{2})+2\beta(4X^{2})$ and
@@
$\begin{split}\frac{d}{dt}(y_{1}-y_{2})^{2}={}&8(y_{1}-y_{2})\bigg{[}(p_{12}-q_{12})(y_{1}-y_{2})(\log\beta)^{\prime}(|y_{1}-y_{2}|^{2})\\ {}&-(p_{12}-q_{12})(y_{2}-y_{1})(\log\beta)^{\prime}(|y_{1}-y_{2}|^{2})\\ {}&+\sum_{j>2}(p_{1j}-q_{1j})(y_{1}-y_{j})(\log\beta)^{\prime}(|y_{1}-y_{j}|^{2})\\ {}&-\sum_{j>2}(p_{2j}-q_{2j})(y_{2}-y_{j})(\log\beta)^{\prime}(|y_{2}-y_{j}|^{2})\bigg{]}.\end{split}$
@@
$\frac{d}{dt}(y_{1}-y_{2})^{2}\leq C(y_{1}-y_{2})^{2},$
and thus $e^{-Ct}(y_{1}-y_{2})^{2}$ is decreasing on $[0,T]$ and initially vanishes. Hence $y_{1}(t)=y_{2}(t)$ on $[0,T]$ and since $T$ was arbitrary this proves $y_{1}(t)=y_{2}(t)$ for all $t$.

### v2

# Stochastic neighborhood embedding and the gradient flow of relative entropy

@@ 1.1. The SNE and t-SNE algorithms
$p_{j|i}=\frac{\exp(-|x_{i}-x_{j}|^{2}/(2\sigma_{i}^{2}))}{\sum_{k\,|\,k\neq i}\exp(-|x_{i}-x_{k}|^{2}/(2\sigma_{i}^{2}))}\quad\textrm{for }i\neq j.$
This represents “the conditional probability that $x_{i}$ would pick $x_{j}$ as its neighbor if neighbors were picked in proportion to their probability density under a Gaussian centered at $x_{i}$[14], where $\sigma_{i}$ is the variance of the Gaussian. The numbers $\sigma_{i}$ are selected according to certain criteria; since this is not the focus of the paper refer the reader to [14]. A word about summation notation is called for. If $i$ is fixed, we use $\sum_{k\,|\,k\neq i}$ to denote the sum over all $k=1,\ldots,n$ with $k$ not equal to $i$. We will use $\sum_{k\neq\ell}$ to denote a double summation over all $k$ and all $\ell$ from $1$ to $n$ with $k\neq\ell$.
We define
@@
(1.2) $\mathcal{C}(Y):=\sum_{\begin{subarray}{c}i\neq j\end{subarray}}p_{ij}\log\frac{p_{ij}}{q_{ij}},$
@@ 1.2. The gradient flow of relative entropy
(1.3) $\begin{split}\frac{dy_{i}}{dt}={}&4\sum_{\begin{subarray}{c}j\,|\,j\neq i\end{subarray}}(p_{ij}-q_{ij})(y_{i}-y_{j})(\log\beta)^{\prime}(|y_{i}-y_{j}|^{2}),\end{split}$
@@ 1.3. Results
    $\sup_{x\in[0,\infty)}(\log\gamma)^{\prime}(x)<\infty.$
@@ 1.4. SNE versus t-SNE
Our results show a marked difference in the behavior of the flow $Y(t)$ in the SNE versus t-SNE cases. Only t-SNE exhibits divergence of points at infinity, and Theorem 1.3 shows that if this occurs one obtains distinct points in the rescaled limit. We believe this is related to the so-called “crowding problem” discussed by van der Maaten and Hinton in [14, Section 3.2]. They note that in SNE, the points $y_{i}$ tend to be pushed towards the center, preventing gaps from occurring between clusters. They give heuristic reasons why the heavier tails of the probability distribution in t-SNE may compensate for this effect, pushing points apart. Indeed, this motivated their construction of the t-SNE algorithm.
The outline of the paper is as follows. In Section 2 we compute the formula (1.3) for the gradient flow. Section 3 is the heart of the paper, in which we prove the main results, Theorems 1.1 to 1.4. In Section 4, we give two further examples. Finally in Section 5, we pose some further questions for future study.
Acknowledgements.  The author is very grateful to Antonio Auffinger for spurring his interest in the mathematical study of t-SNE, and for some helpful comments. The author also thanks the referees for suggesting some improvements to the exposition.
@@ 2. The gradient flow of relative entropy
In this section we prove the formula (1.3) for the gradient flow of the entropy $\mathcal{C}$. The gradient of $\mathcal{C}$ was calculated for $\beta=e^{-x}$ and $\beta(x)=(1+x)^{-1}$ in [7, 14] (see also [4, Chapter 16]) and its extension to general $\beta$ is straightforward. However, we include the details here for the convenience of the reader [^1] The author thanks one of the referees for pointing out a simplification of his original argument.. Recalling (1.1), we write
(2.1) $q_{ij}=\frac{\beta(|y_{i}-y_{j}|^{2})}{Z},\quad\textrm{for }Z:=\sum_{i\neq j}\beta(|y_{j}-y_{j}|^{2}).$
We have
$\begin{split}\nabla_{y_{m}}\mathcal{C}={}&-\sum_{i\neq j}p_{ij}\nabla_{y_{m}}\log q_{ij}\\ ={}&-\sum_{i\neq j}p_{ij}\left(\nabla_{y_{m}}\log\beta(|y_{i}-y_{j}|^{2})-\nabla_{y_{m}}\log Z\right)\\ ={}&-\sum_{i\neq j}p_{ij}(\log\beta)^{\prime}(|y_{i}-y_{j}|^{2})\nabla_{y_{m}}|y_{i}-y_{j}|^{2}+\frac{1}{Z}\nabla_{y_{m}}\sum_{i\neq j}\beta(|y_{i}-y_{j}|^{2}),\end{split}$
where for the last line we used the fact that $\sum_{i\neq j}p_{ij}=1$. Observe that
$\nabla_{y_{m}}|y_{i}-y_{j}|^{2}=2\delta_{mi}(y_{m}-y_{j})+2\delta_{mj}(y_{m}-y_{i}).$
Hence
$\begin{split}\nabla_{y_{m}}\mathcal{C}={}&-4\sum_{j\,|\,j\neq m}p_{mj}(y_{m}-y_{j})(\log\beta)^{\prime}(|y_{m}-y_{j}|^{2})+\frac{4}{Z}\sum_{j\,|\,j\neq m}(y_{m}-y_{j})\beta^{\prime}(|y_{m}-y_{j}|^{2})\\ ={}&-4\sum_{j\,|\,j\neq m}(p_{mj}-\frac{\beta(|y_{m}-y_{j}|^{2})}{Z})(y_{m}-y_{j})(\log\beta)^{\prime}(|y_{m}-y_{j}|^{2})\\ ={}&-4\sum_{j\,|\,j\neq m}(p_{mj}-q_{mj})(y_{m}-y_{j})(\log\beta)^{\prime}(|y_{m}-y_{j}|^{2}),\end{split}$
@@ 3. Proofs of the main results
$\mathcal{C}(Y)=\sum_{i\neq j}p_{ij}\log\frac{p_{ij}}{q_{ij}},$
we have $\mathcal{C}(Y(t))\leq\mathcal{C}(Y_{0})$ and hence,
$-\sum_{i\neq j}p_{ij}\log q_{ij}\leq-\sum_{i\neq j}p_{ij}\log p_{ij}+\mathcal{C}(Y_{0})\leq C,$
for a uniform constant $C$. Since each term $-p_{ij}\log q_{ij}$ in the sum is positive, we have for $i\neq j$,
$-p_{ij}\log q_{ij}\leq C.$
This implies in particular that
$\log q_{12}\geq-C/p_{12}=:-C^{\prime}$
for a uniform $C^{\prime}$. Here we recall our assumption that each $p_{ij}$ (for $i\neq j$) is strictly positive and our uniform constants may depend on the $p_{ij}$. Exponentiating gives
@@
for a uniform positive $c^{\prime}$.
@@
$\begin{split}\frac{d}{dt}\sum_{i=1}^{n}|y_{i}|^{2}={}&8\sum_{i\neq j}(p_{ij}-q_{ij})y_{i}\cdot(y_{i}-y_{j})(\log\beta)^{\prime}(|y_{i}-y_{j}|^{2})\\ ={}&8\sum_{i<j}(p_{ij}-q_{ij})y_{i}\cdot(y_{i}-y_{j})(\log\beta)^{\prime}(|y_{i}-y_{j}|^{2})\\ {}&+8\sum_{i<j}(p_{ij}-q_{ij})y_{j}\cdot(y_{j}-y_{i})(\log\beta)^{\prime}(|y_{i}-y_{j}|^{2})\\ ={}&8\sum_{i<j}(p_{ij}-q_{ij})|y_{i}-y_{j}|^{2}(\log\beta)^{\prime}(|y_{i}-y_{j}|^{2}),\end{split}$
where $\sum_{i<j}$ denotes the double summation of $i$ and $j$ with $i<j$. Now for $i\neq j$, write
@@
$\begin{split}\frac{d}{dt}S={}&\frac{8}{Z}\left(p_{12}Z-A_{12}\right)\log\frac{A_{12}}{Z}+\frac{8}{Z}\sum_{\begin{subarray}{c}i<j\,|\,(i,j)\neq(1,2)\end{subarray}}\left(p_{ij}Z-A_{ij}\right)\log\frac{A_{ij}}{Z}\\ \leq{}&-\frac{4c^{\prime}}{Z}p_{12}ZS+\frac{8}{Z}\sum_{\begin{subarray}{c}i<j\,|\,(i,j)\neq(1,2),\\ p_{ij}Z-A_{ij}<0\end{subarray}}\left(p_{ij}Z-A_{ij}\right)\log\frac{A_{ij}}{Z},\end{split}$
@@
$\begin{split}\frac{d}{dt}S={}&-cS+\frac{8}{Z}\sum_{\begin{subarray}{c}i<j\,|\,(i,j)\neq(1,2),\\ p_{ij}Z-A_{ij}<0\end{subarray}}\left(p_{ij}Z-A_{ij}\right)\log p_{ij}\\ \leq{}&-cS+C<0,\end{split}$
@@
We provide an example with three points in $\mathbb{R}$. Write $p_{12}=p_{23}=a$ so that $p_{13}=1/2(1-4a)$ with $a\in(0,1/4)$ to be determined. Assume that $y_{1}(t)=X(t)$, $y_{2}(t)=0$ and $y_{3}(t)=-X(t)$ so that the points are symmetric about the origin, with $X(0)>0$. Recalling (2.1), we have $Z=4\beta(X^{2})+2\beta(4X^{2})$ and
@@
$\begin{split}\frac{d}{dt}|y_{1}-y_{2}|^{2}={}&8(y_{1}-y_{2})\cdot\bigg{[}(p_{12}-q_{12})(y_{1}-y_{2})(\log\beta)^{\prime}(|y_{1}-y_{2}|^{2})\\ {}&-(p_{12}-q_{12})(y_{2}-y_{1})(\log\beta)^{\prime}(|y_{1}-y_{2}|^{2})\\ {}&+\sum_{j>2}(p_{1j}-q_{1j})(y_{1}-y_{j})(\log\beta)^{\prime}(|y_{1}-y_{j}|^{2})\\ {}&-\sum_{j>2}(p_{2j}-q_{2j})(y_{2}-y_{j})(\log\beta)^{\prime}(|y_{2}-y_{j}|^{2})\bigg{]}.\end{split}$
@@
$\frac{d}{dt}|y_{1}-y_{2}|^{2}\leq C|y_{1}-y_{2}|^{2},$
and thus $e^{-Ct}|y_{1}-y_{2}|^{2}$ is decreasing on $[0,T]$ and initially vanishes. This implies that $|y_{1}-y_{2}|^{2}$ remains zero for $t\in[0,T]$. Hence $y_{1}(t)=y_{2}(t)$ on $[0,T]$ and since $T$ was arbitrary this proves $y_{1}(t)=y_{2}(t)$ for all $t$.
@@ 4. Examples
5. Further questions
We are still at the early stages of understanding the behavior of solutions $Y(t)$ of the gradient flow of relative entropy. Some follow-up questions include:
  1. (i)
  2. (ii)
  3. (iii)
  4. (iv)
  5. (v)
    If $Y(t)$ converges to a critical point as $t\rightarrow\infty$, what is the rate of convergence? (This is of practical importance for the users of the t-SNE algorithm).